OpenStudy (anonymous):

simplify the following expression log(25x^3)+3log(1/x)

OpenStudy (caishaax3):

exact form or decimal?

OpenStudy (caishaax3):

exact form: log(25) decimal form: 1.39794000...

OpenStudy (anonymous):

\[\log 25x^3{} + 3\log\frac{ 1 }{ x }=\log25 + 3\log x + 3\log x^{-1} = \log 25 + 3.\log x - 3.\log x = \log 25 = 2.\log 5\]